OpenStudy (anonymous):

Ok guys next and last question 5(x+3) - ( 5x+15)=0

OpenStudy (anonymous):

5(x+3) - ( 5x+15)=0 5(x+3) = 5x+15 5x+15 = 5x+15 5x = 5x x = x x= 1

OpenStudy (anonymous):

Ah I can see whe I messed up. Thank you

OpenStudy (anonymous):

no prob

OpenStudy (cggurumanjunath):